# Paramanu-Ayn: An Efficient Novel Generative and Instruction-tuned Language Model for Indian Legal Case Documents

### v1

@@ Abstract
In this paper, we present Paramanu-Ayn, a language model based exclusively on case documents of the Supreme Court of India, the Constitution of India, and the Indian Penal Code. The novel Auto Regressive (AR) decoder based model is pretrained from scratch at a context size of 8192. We evaluated our pretrained legal model on perplexity metrics. We also instruction-tuned our pretrained model on a set of 10,763 instructions covering various legal tasks such as legal reasoning, judgement explanation, legal clause generation, legal drafting, legal contract drafting, case summarization, constitutional question-answering, etc. We also evaluated the responses of prompts for instruction-tuned models by GPT-3.5-Turbo on clarity, relevance, completeness, and legal reasoning metrics in a scale of 10. Our model can be run on CPU and achieved 42.46 tokens/sec CPU inference speed. We found that our models, despite not being pretrained on legal books, various legal contracts, and legal documents, were able to learn the domain knowledge required for drafting various legal contracts and legal clauses, and generalize to draft legal contracts and legal clauses with limited instruction tuning. Hence, we conclude that for a strong domain-specialized generative language model (such as legal), very large amounts of data are not required to develop models from scratch. We believe that this work is the first attempt to make a dedicated generative legal language model from scratch for Indian Supreme Court jurisdiction or in legal NLP overall. We plan to release our Paramanu-Ayn model at https://www.bharatgpts.com.
@@ 1 Introduction
Legal text is characterized by its unique syntax and specialized vocabulary. This poses a distinct linguistic challenge for language models ((Chalkidis et al., 2020), (Niklaus and Giofré, 2022)). Pretrained Large Language Models (LLMs) (LLaMa (Touvron et al., 2023), InLegalBERT (Paul et al., 2023), Longformer (Beltagy et al., 2020)) are often trained on domain specific data for domain adaptation (Yao et al., 2021), (Gururangan et al., 2020). Instead of following the domain adaptation method of LLMs for legal language modeling for legal documents, we focused to pretrain from scratch a generative legal language model for case documents of Supreme Court of India. This avoids high inference latency, high cost of training, and non-specialised tokenizer issue of existing LLMs. Instead of developing humongous models that are believed to be good at all domains and for all tasks, we believe in distributed specialized domain-expert machines. Hence, we want to develop domain-specific smaller models from scratch for resource-constrained computing setting.
Our legal model is based on Paramanu (Niyogi and Bhattacharya, 2024), released earlier by us. We have trained an auto-regressive models from scratch at a context size of 8192 on a single NVidia A100-PCIE-40GB GPU. Our work is an attempt to make dedicated domain specialized models from scratch rather than performing continual pretraining of existing LLMs for domain adaptation. Our models are much smaller in size having only a few million parameters. Hence, our models are very fast in inference without requiring any quantization of weights, and our legal models can be run on CPU without need of GPU.
Our main contributions are as follows:
@@
  • We developed first ever dedicated AR legal model from scratch for Indian Supreme Court at context size of 8192 respectively on a single GPU.
  • We evaluated our legal pretrained models on validation perplexity, and on model FLOPs Utilization (MFU) metric for pretraining. Table 1 shows the validation perplexity and MFU metrics of pretraining.
  • We also instruction tuned our Paramanu-Ayn legal pretrained model on 10,763 legal instructions and demonstrated various legal tasks capabilities of our model.
  • We also evaluated the responses of Paramanu-Ayn-instruct by GPT-3.5-Turbo on clarity, relevance, completeness, and legal reasoning metrics in a scale of 10. Table 3 and Table 4 show the respective evaluation of GPT-3.5-Turbo related to general legal instructions and legal instructions related to Supreme Court cases in India.
@@ 2 Related Work
Talking about legal evaluation, of law-specific LLMs needs benchmarks that focus on legal tasks. Since legal laws and processes vary by country and country specific, developing a single global legal LLM evaluation benchmark is not feasible to our opinion.
@@ 3.1 Language Modeling
This objective of the language modeling can formally described as maximizing the probability of a sequence of tokens $w1,w2,\dots,w_{N}$
$P(w_{1},w_{2},\ldots,w_{n})=\prod_{i=1}^{n}P(w_{i}\,|\,w_{1},w_{2},\ldots,w_{i-1})$ (1)
@@
The performance of a language model is generally being evaluated using the total cross-entropy loss, i.e, the negative log-likelihood of the observed data under the model under consideration, which for a given dataset is defined as:
Avg $\text{Avg Loss}=-\frac{1}{N}\sum_{i=1}^{n}\log(P(w_{i}\,|\,w_{1},w_{2},\ldots,w_{i-1}))$ (2)
Lower the loss better is the model but just computing the loss may be not intuitive. Therefore, perplexity is a metric to evaluate the performance of a given language model which is the exponent of the average loss.
@@ 3.2 Model flops utilization (MFU)
Model Flops Uitilization (MFU) (et al., 2022) (Korthikanti et al., 2022) estimate is the ratio of the observed throughput (tokens-per-second), relative to the theoretical maximum throughput of a system at peak FLOPs. Model flops utilization (MFU) estimate the number of flops we do per iteration. It quantifies how efficiently the GPUs are utilized in model training.
@@ 4.1 Pretraining Data
The pretraining data is an extension version of ILDC (Malik et al., 2021) dataset. We have scraped more Supreme Court of India court case documents till December 2023. Apart from Supreme Court cases, we also have the Constitution of India in English and the Indian Penal Code, which we scraped using public sources.
@@ 4.2 Instruction Tuning Data
The instruction dataset is a cumulated corpus of various existing publicly available dataset (NebulaSense, 2023; Ayansk11, 2023). This results in a total of 12,270 instructions, and after removing duplicate instructions, it resulted in around 10,763 instructions covering various legal tasks such as legal memo drafting, legal petition drafting, legal contract generation such as employee contract drafting, NDA drafting, legal clause generation, legal argument challenging, legal advice based on an outcome of a Supreme Court case, legal case summarization and key findings explanation, legal precedents identification, legal question answering, etc. Some of the tasks in the instruction dataset are mentioned in Section A in the appendix.
@@ 5 Training
For pretraining Paramanu-Ayn (8192 context size), we performed 95%-5% data split for pretraining, as we wanted to use most of the dataset for pretraining. We reported the validation perplexity of our pre-trained model.
For instruction-tuning, we split the dataset into 90%-10% training and testing sets. Since our models are smaller in size, we performed supervised full fine-tuning of Paramanu-Ayn for 3 epochs on the accumulated 10,763 instructions and reported both training and validation loss for instruction tuning using cosine, constant, and linear learning rate scheduler. We found that cosine learning rate scheduler results in lowest validation loss for our models.
Model Perplexity MFU
Paramanu-Ayn (8192) 1.4582095 41.34760
Table 1: Perplexity and MFU metrics of Paramanu-Ayn pretrained models.
Model Inference Speed (tokens/sec)
Paramanu-Ayn 42.4575
Table 2: CPU (AMD epyc 7252 8-core processor, RAM:128GB RAM) inference speed of Paramanu-Ayn models.

### v2

# Paramanu-Ayn: Pretrain from scratch or Continual Pretraining of LLMs for Legal Domain Adaptation?

@@ Abstract
In this paper, we present Paramanu-Ayn, a collection of language models based exclusively on legal case documents of India. In particular, the court case documents of the Supreme Court of India (SCI), and text of the Constitution of India and the Indian Penal Code are utilized. The novel Auto-Regressive (AR) decoder-only language model is pretrained from scratch at a context size of 8192 on a single A100-PCIE-40GB GPU with an efficient MFU (Model FLOPs Utilization) of 41.35. Our model size is only 97 million parameters. We also trained a tokenizer specialized for legal corpora based on BPE. We evaluated our pretrained legal model on perplexity metric, and did zero-shot evaluation on two tasks: case judgement prediction with explanation and abstractive legal case summarization. Our pretrained legal model outperformed LlaMa-2 7B and Gemini-Pro on zero-shot test accuracy for the legal case judgement prediction with explanation task by close to 2% points, despite being smaller in size by close to 72 times compared to these 7B LLMs. We also reported automatic lexical metrics of zero-shot abstractive summarization generation of our Paramanu-Ayn model across varying summary lengths and compared with decoder-only LLMs at fixed summary length generation (5000 tokens) using ROUGE-1, ROUGE-L, BLEU, METEOR, and BERTScore metrics. Paramanu-Ayn outperformed the models in both BLEU and METEOR metrics by more than 10% percentage points, and also on BERTscore by close to 4% points. Furthermore, we evaluated our model on popular LLM commonsense and mathematical benchmarks for a holistic evaluation despite the model being pretrained exclusively on legal court cases documents. Paramanu-Ayn scored impressively on all the benchmarks. Paramanu-Ayn outperformed Llama-1, Llama-2, and Falcon LLMs on mathematical benchmarks (AGIEVAL-AQuA-RAT, AGIEVAL-SAT-Math) at zero-shot setting. We also instruction-tuned our pretrained model on a set of 10,763 instructions covering various legal tasks such as legal reasoning, judgement explanation, legal clause generation, legal drafting, legal contract drafting, case summarization, constitutional question-answering, etc. We also evaluated the responses of prompts for instruction-tuned models for both legal instructions related to Supreme Court of India cases and on drafting legal clauses and legal contracts by GPT-3.5-Turbo on clarity, relevance, completeness, and legal reasoning metrics. Paramanu-Ayn-instruct model scored more than 8 (on a scale of 10) on all the metrics, namely, clarity, relevance, completeness, and legal reasoning. Our model can be run on CPU and achieved 42.46 tokens/sec CPU inference speed. We found that our models, despite being pretrained exclusively only on court case documents, were able to learn the domain knowledge required for drafting various legal contracts and legal clauses, and generalize to draft legal contracts and legal clauses with limited instruction tuning. Hence, we conclude that for a strong domain-specialized generative language model (such as legal), small domain specialized pretraining from scratch is more cost effective. Our 97 million parameters model was pretrained from scratch for only 185 A100 training hours but is still equally competitive or even better than adapting LLMs for legal domain tasks. We believe that this work is the first attempt to make an exclusive generative legal language model from scratch for Indian Supreme Court jurisdiction or in legal NLP overall, and offer an environmental friendly cost efficient approach for legal domain adaptation.
@@ 1 Introduction
Legal text is characterized by its unique syntax and specialized vocabulary. This poses a distinct linguistic challenge for language models (Chalkidis et al., 2020; Niklaus and Giofre, 2023). Pretrained Large Language Models (LLMs) such as LLaMa (Touvron et al., 2023), Longformer (Beltagy et al., 2020), etc. are often trained on domain specific data for domain adaptation (Yao et al., 2021; Gururangan et al., 2020; Colombo et al., 2024). We propose a different approach. Instead of following the domain adaptation method of LLMs for legal language modeling for legal documents, we focused to pretrain from scratch a generative legal language model for case documents of Supreme Court of India. This avoids high inference latency, high cost of training, and non-specialised tokenizer and the misalignment of domain specialised tokenizers and embeddings with the existing embeddings of large language models (LLMs) via continual pretraining with vocabulary expansion of the existing LLMs tokenizers. Instead of developing humongous models that are believed to be good for all domains and all tasks, we believe in distributed specialized domain-expert machines. Hence, we want to develop domain-specific smaller models from scratch for resource-constrained computing setting.
Our model is based on the Transformer Decoder architecture Vaswani et al. (2017). We have trained a collection of an auto-regressive generative legal language models from scratch at a context size of 8192 on a single NVidia A100-PCIE-40GB GPU. Our work is an attempt to make dedicated domain specialized models from scratch rather than performing continual pretraining of existing LLMs for domain adaptation. Our models are much smaller in size and have just 97 million parameters. Hence, our models are very fast in inference without requiring any quantization of weights. Our legal models can be run on CPUs or even in smartphones without the need of GPUs.
Our main contributions in this work are as follows:
@@
  • We developed the first ever dedicated collection of generative legal AR models pretrained from scratch for Indian Supreme Court at context size of 4096 and 8192 respectively on a single GPU.
  • We evaluated our legal pretrained models on validation perplexity, and on model FLOPs Utilization (MFU) metric for pretraining. Table 2 shows the metrics.
  • We performed zero-shot evaluation of our legal model on the case judgement prediction with explanation task. Table 3 shows that our model outperformed LLMs such as LLaMa-2 7B and Gemini-Pro on zero-shot evaluation.
  • We performed automatic lexical evaluation of zero-shot summary generation of our legal model and decoder-only LLMs on the case summarization task. Table 4 shows the results of the evaluation of Paramanu-Ayn across abstractive summary generation across varying lengths of token generation. Table 5 compares the automatic lexical metrics of zero-shot evaluation of LLMs versus Paramanu-Ayn on zero-shot case abstractive summarization task.
  • We also evaluated Paramanu-Ayn on zero-shot evaluation on common LLM benchmarks across commonsense reasoning, logical, analytical, and mathematical reasoning. Table 6 shows the benchmark evaluation of Paramanu-Ayn on these tasks. On the mathematical benchmark, AGIEVAL-AQuA-RAT, and AGIEVAL-SAT-Math, Paramanu-Ayn 97M outperformed various LLMs like Falcon (7B,40B), Llama-1 (7B, 13B, 33B), Llama-2 (13B, 34B) respectively. For abstract reasoning on AGIEVAL-LSAT-AR, Paramanu-Ayn 97M also outperformed Falcon and Llama LLMs.
  • We further instruction-tuned our Paramanu-Ayn legal pretrained model on 10,763 legal instructions and demonstrated various legal task capabilities.
  • We evaluated the responses of Paramanu-Ayn-instruct by GPT-3.5-Turbo on clarity, relevance, completeness, and legal reasoning metrics. Table 8 shows the results.
@@ 2 Related Work
Talking about legal evaluation, of law-specific LLMs needs benchmarks that focus on legal tasks. Since legal laws and processes vary by country and country specific, developing a single global legal LLM evaluation benchmark may not be desirable in our opinion.
@@ 3.1 Language Modeling
This objective of the language modeling can be formally described as maximizing the probability of a sequence of tokens $w_{1},w_{2},\dots,w_{N}$
$P(w_{1},w_{2},\ldots,w_{n})=\prod_{i=1}^{n}P(w_{i}\,|\,w_{1},w_{2},\ldots,w_{i-1})$
@@
The performance of a language model is generally evaluated using the total cross-entropy loss, i.e, the negative log-likelihood of the observed data under the model under consideration, which for a given dataset is defined as:
$Loss=-\frac{1}{N}\sum_{i=1}^{n}\log(P(w_{i}\,|\,w_{1},w_{2},\ldots,w_{i-1}))$
Lower the loss better is the model; however, just computing the loss may not be intuitive. Therefore, Perplexity is a metric to evaluate the performance of a given language model which is the exponent of the average loss.
$Perplexity=\exp\left(Loss\right)$
3.2 Rotary Position Embedding (RoPE)
Transformer-based models rely on positional embeddings to encode position and relative location information of words in a text. Rotary Position Embedding (RoPE) is a position encoding technique proposed by (Black et al., 2022). Instead of adding positional embeddings or relative positional embeddings to token embeddings, RoPE rotates the token embedding by a fixed factor ($\theta$) in the higher-dimensional space to encode relative positional embeddings. In other words, RoPE encodes the absolute positions with a rotation matrix and meanwhile incorporates the explicit relative position dependency in self-attention formulation. The intuition behind RoPE is that we can represent the token embeddings as complex numbers and their positions as pure rotations that we apply to them. By exploiting the nature of rotations, the dot product used in self-attention will have the property for preserving relative positional information while discarding absolute position.
3.3 Root Mean Square Normalization (RMSNorm)
To improve the training stability, some LLMs (Chinchilla (Hoffmann et al., 2022), LLaMa (Touvron et al., 2023)) have normalized the input of each transformer sub-layer, instead of normalizing the output using RMSNorm normalizing function as introduced by (Zhang and Sennrich, 2019). RMSNorm normalizes the activations based on their root mean square (RMS) value instead of normalizing the inputs based on their mean and variance. RMSNorm accelerates the training and inference with similar performance in these large models. It is reported that replacing LayerNorm (Ba et al., 2016) with RMSNorm can achieve comparable performance and improve training and inference time by 7-64%. Narang et al. (2021) showed that RMSNorm improves the pre-training speed by 5% compared with the LayerNorm baseline.
@@ 3.4 Model FLOPs Utilization (MFU)
Model FLOPs Utilization (MFU) (Chowdhery et al., 2024) estimate is the ratio of the observed throughput (tokens-per-second) relative to the theoretical maximum throughput of a system at peak FLOPs. Model flops utilization (MFU) estimate the number of flops (floating point operations) done per iteration. It quantifies how efficiently the GPUs are utilized in model training.
3.5 Maximal Update Parameterization
As the size of large language models (LLMs) and the scale of the dataset used in pretraining are expensively large, it is not feasible to perform hyperparameter tuning in LLMs. Yang et al. (2021) used a technique called maximal update parameterization ($\mu P$) to transfer the hyperparameters learnt from tuning of a small model to a larger model and found that the optimal hyperparameter values become stable across neural network sizes when the models have been parameterized using ($\mu P$).
@@ 4.1 Pretraining Data
The pretraining data is an extended version of ILDC (Malik et al., 2021) dataset which initially contained 34,816 Supreme Court of India cases. We extracted all the publicly available case proceedings from May 2020 to December 2023 (3,046 case documents) and augmented the dataset. Apart from Supreme Court cases, we also have added the Constitution of India (in English) and the Indian Penal Code (in English), which we scraped using public sources. Table 1 displays the statistics for our clean preprocessed pretraining corpus. As the number of tokens vary due to the variation of tokenizer, we reported the total number of (whitespace-separated) words instead.
4.2 Data Cleaning and Preprocessing
Case proceedings are unstructured documents and have different formats and sizes. In addition, there are typos and spelling mistakes (since these are typed during the court hearing), thereby making it challenging to preprocess. We used regular expressions to remove the noisy text (unreadable characters, special characters, unnecessary whitespaces, etc.) and meta-information (e.g., initial portions of the document containing case number, judge name, dates, and other meta information) from the proceedings. In Supreme Court case proceedings, the decisions are written towards the end of the document. These end section(s) directly stating the decision have been kept instead of removing from the documents like the ILDC dataset.
For both the Constitution of India and the Indian Penal Code, we used the tesseract tool to extract the text from the pdf files and also removed the first few pages of the pdf to remove title page, acknowledgments, table of contents, and other irrelevant information.
Pre-training Corpus Timeline #Words
38,222 SCI Cases 1947 – 2023 142,374,414
Constitution of India 2024 139,956
Penal Code of India 2024 77,633
Total 1947 – 2024 142,592,003
Table 1: Pretraining Corpus
@@ 4.3 Instruction Tuning Data
The instruction dataset is a cumulative corpus of various existing publicly available datasets (NebulaSense, 2023; Ayansk11, 2023). This results in a total of 12,270 instructions; after removing duplicates, it resulted in 10,763 instructions covering various legal tasks such as legal memo drafting, legal petition drafting, legal contract generation such as employee contract drafting, NDA drafting, legal clause generation, legal argument challenging, legal advice based on an outcome of a Supreme Court case, legal case summarization and key findings explanation, legal precedents identification, legal question answering, etc.
5 Tokenization
We trained Byte-Pair encoding (BPE) (Sennrich et al., 2016) algorithm using Sentencepiece (Kudo and Richardson, 2018) module on the pretraining data from scratch to develop legal specialised tokenizer that aims to learn the intricacies of legal terminology. During pre-tokenization, NFC normalization was performed on the processed data; digits are split into individual tokens and unknown UTF-8 characters were reduced to byte granularity for improving the arithmetic learning ability of the pretrained model. In NFC, characters are normalized by composing them into their canonical form. This means that composed characters (like “é”) are represented as a single code point rather than as a combination of base characters and diacritical marks (like ‘e’ followed by an accent mark). This helps in ensuring consistency and proper comparison of text, as different sequences of characters that visually represent the same text can be treated as identical. We treat the data as a sequence of bytes rather than Unicode characters, and we include each of the 256 bytes as tokens too. We removed the overlapping tokens from the trained tokenizer by removing the duplicated tokens in the tokenizer to develop our tokenizer legal specialised, compact, optimized, and effective. The size of the final tokenizer vocabulary is 15,575.
6 Model Architecture
The model architecture of Paramanu-Ayn uses a scaled version of RoPE embedding (Su et al., 2024) with $\theta$=10,000. We scaled the RoPE embedding through a shrinking factor by dividing the target context length $y$ by $permissible\_context\_size\_length$ on single GPU, keeping all other hyperparameters fixed such as batch size, vocabulary size, etc. This allows every $position\_ids$ to be divided by the shrinking ratio in the RoPE embedding methodology. For instance, if the $permissible\_context\_size\_length$ for a given physical memory hardware is 256, then we apply shrinking factor of 16 for target context size of 4096 on Single A100 40G chip during pretraining. Then, a token with $position\_ids$ = 4000 becomes 4000/16 = 250, and the neighbouring token 4001 becomes 4001/16 = 250.06, to be within 0 to 256. This is how we can capture higher context size during pretraining on limited physical memory required to pretrain model at higher context size outside the $permissible\_context\_size\_length$. This modification allows us to pretrain models from scratch at much higher context size than the physical memory required for pretraining. Hence, with limited physical memory and limited GPUs, we can pretrain generative language models from scratch at much higher desired context size. Applying this technique allowed us to pretrain our legal model at a context size of 8192 on a single NVIDIA A100-PCIE-40GB GPU. The shrinking factor was set to 32 to achieve a target context size of 8192.
The model architecture of Paramanu-Ayn uses RMSNorm (Zhang and Sennrich, 2019) as a pre-normalizaion layer with norm_epsilon = 1e-5, uses SwiGLU (Shazeer, 2020) an activation function for non-linearity by replacing the standard ReLU activation function. The dimension of Paramanu-Ayn is 768 with 12 layers with a feedforward layer hidden dimension of 2048. Following (Su et al., 2024), we remove all biases from dense layers to improve the training stability of Paramanu-Ayn. The model Paramanu-Ayn uses grouped-query attention (GQA) (Ainslie et al., 2023) to reduce inference-time memory footprint and accelerate inference at run time. GQA divides query heads into $G$ groups, each of which shares a single key head and value head. GQA-G refers to grouped-query with $G$ groups. In GQA, we divide the $N$ total self-attention heads into groups, where key and value heads are shared within each group; GQA-1, with a single group and therefore single key and value head, is equivalent to multi-query attention (MQA) (Shazeer, 2019), while GQA-H, with groups equal to number of heads, is equivalent to multi-head attention (MHA). We also used weight tying (Press and Wolf, 2017) to improve the performance of language models by sharing the weights of the embedding and softmax layers.
@@ 7 Training
For pretraining Paramanu-Ayn (8192 context size), we performed 95%-5% data split for pretraining, as we wanted to use most of the dataset for pretraining. We reported the validation perplexity of our pre-trained model. Figure 1 shows the training perplexity curve against the number of tokens.
Figure 1: Training Perplexity versus number of Tokens in billions (represented in G, 1G=1 billion)
[…]
$lr_{decay\_ratio}=\frac{t-warmup_{steps}}{lr_{decay\_steps}-warmup_{steps}}$
where $t$ is the current training step. We set the maximum learning rate ($lr$) to 0.003 (max), weight decay to 0.1. Paramanu-Ayn is pretrained on around a total of 26 billion tokens. Fig 1 shows the plot of training perplexity with respect to every billion tokens. We trained with fused AdamW optimizer with $\beta_{1}$ = 0.9, $\beta_{2}$ = 0.95, a gradient clipping of 1.0, and weight decay of 0.1. To further speedup training, we used BF16 mixed precision training. Paramanu-Ayn was trained for 185 A100 hours. For our experiments and modeling, we implemented our code using Pytorch 2.0, in-house optmized CUDA kernels and used torch.compile feature for every model.
For instruction-tuning, we split the dataset into 90%-10% training and testing sets. Since our models are smaller in size, we performed supervised full fine-tuning of Paramanu-Ayn for 3 epochs on the accumulated 10,763 instructions and trained in different setting of experiments using cosine, constant, and linear learning rate scheduler with ($lr$) set to $2$e55-5- 5, gradient clipping of 1.0, warmup ratio of 0.05 and no weight decay. We found that cosine learning rate scheduler results in the lowest validation loss for our models.
8 Carbon Footprint
To measure carbon footprint for our pretraining, we follow Touvron et al. (2023):
$tCO_{2}eq=MWh\times 0.385$
The power consumption can be calculated as
[…]
where PUE is Power Usage Effectiveness.
We observed during pretraining that our single A100 40G consumes 250 Watt consistently. Therefore, our pretraining process consumed 50.875 kW, resulting in only 0.0196 tCO2eq respectively. This makes our model way more environmental-friendly as compared to pretraining/continual pretraining of LLMs.
9 Evaluation
We evaluated Paramanu-Ayn and LLMs including (a) LlaMa-3 8B, (b) Mixtral 7B, (c) Gemini Pro on (1) Legal Judgement Case Prediction task (Nigam et al., 2024) using zero-shot evaluation which is exclusively for Supreme Court of India cases, (2) Supreme Court of India abstractive summarization (In-Abs) (Shukla et al., 2022) test dataset of 100 case documents using zero-shot evaluation on summary generation of 5000 tokens. We further studied the long summary generation capability of Paramanu-Ayn by generating summaries of varying lengths of 1024, 2048, 4096, 5000, 6000, and 8192 tokens and compared the lexical evaluation metrics like ROUGE-1, ROUGE-L, METEOR, BLEU, and semantic evaluation metrics like BERTScore. ROUGE family of metrics measure the textual overlap (uni-grams) between the model-generated summaries and the reference summaries. METEOR calculates the harmonic mean of unigram precision and recall. For summary evaluation, METEOR is used to calculate the unigram overlap between a model-generated summary and the gold standard summary. BERTScore uses BERT to compute the similarity scores between the token level embeddings of the model-generated and reference summaries. We also evaluated Paramanu-Ayn on zero-shot evaluation on MMLU-legal tasks (Professional Law, Jurisprudence, International Law) and analytical reasoning such as AGIEVAL-LSAT-LR (Law school admission tests) which measure the reasoning and analytical skills of prospective law students. These tests include sections on logical reasoning, reading comprehension, and analytical reasoning. Furthermore, we went ahead to explore Paramanu-Ayn capabilities on various commonsense reasoning benchmarks such as ARC-easy, ARC-challenge, SciQ, PIQA, WSC, COPA, mathematical reasoning such as AGIEVAL-AQuA-RAT (GRE, GMAT multiple-choice math questions), AGIEVAL-SAT-Math (SAT multiple-choice math questions) (Zhong et al., 2024), AGIEVAL-MATH (Hendrycks et al., 2021b) (formal logical reasoning such as MMLU-Formal-Logic, causal reasoning task like COPA (selecting a conclusion causally connected to a premise), language coreference resolution task such as WSC (Levesque et al., 2012), classification task such as WIC (Pilehvar and Camacho-Collados, 2019) and natural language inference benchmark such as QNLI (Wang et al., 2018). We also reported the scores on zero-shot evaluation on MMLU-STEM, MMLU-Econometrics, and across all MMLU tasks.
For Paramanu-Ayn-instruct, we evaluated the responses of the test instructions-tuning dataset using GPT-3.5-Turbo on clarity, relevance, completeness, and legal reasoning in a scale of 10.